OpenStudy (anonymous):

X+3/t=r2 solve for x HELP!

OpenStudy (solomonzelman):

x+3 is the numerator.

OpenStudy (solomonzelman):

right ?

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{x+3}{t}=t^2 }\) like this ?

OpenStudy (solomonzelman):

if this is the question, then 1) multiply both sides times 3, 2) subtract 3 from both sides.